Question involvin radicals: parentheses, 6*sq.root of 2 minus 2* sq. root of 6, parentheses, everything squared.
Answer by mukhopadhyay(490)

[6sqrt(2) - 2sqrt(6)]^2
= [6sqrt(2)]^2 - 2*6*sqrt(2)*2*sqrt(6) + [2sqrt(6)]^2
= 36*2 - 24 sqrt(12) + 4*6
= 72 - 48 sqrt(3) + 24
= 96 - 48 sqrt(3)
= 48[2 - sqrt(3)]